ROTATIONAL SPECTRA, STRUCTURES AND MODELING OF THE SO_{2}–OCS, SO_{2}–CS_{2} AND SO_{2}–CO_{2} COMPLEXES

\underline{SEAN A. PEEBLES}, LINGHONG SUN, IOANNIS I. IOANNOU and ROBERT L. KUCZKOWSKI, Department of Chemistry, University of Michigan, 930 N. University Ave., Ann Arbor, MI 48109-1055.

The rotational spectra of the SO_{2}–OCS and SO_{2}–CS_{2} weakly bound dimers have been assigned by pulsed-nozzle Fouriertransform microwave spectroscopy. Both complexes have C_{S} symmetry with the SO_{2} straddling the linear monomer. The SO_{2} axis makes an angle with the linear monomer of 18.3^{∘} and 9.8^{∘} respectively in the two complexes. This is in contrast to the structure of SO_{2}–CO_{2}.^{a} This complex has effective C_{2v} symmetry with the SO_{2} axis perpendicular to the CO_{2}. The structural features and dipole moments of these three species will be discussed and compared with results obtained from a semi-empirical model employing electrostatic, dispersion and repulsion interactions.

aL.H. Sun, I.I. Ioannou and R.L. Kuczkowski, Mol. Phys., 88, (1996), 255